A large garage with well-insulated walls and containing 700 m3 of air at 305 K is heated at constant pressure (atmospheric). Consider air to be an ideal diatomic gas. (a) Determine the energy (in kJ) required to increase the temperature of the air in the building by 1.70°C. kJ (b) Determine the mass (in kg) this amount of energy could lift through a height 1.50 m. kg
Answer #1

Solution)

Part a)

Here, Energy required, Q=7/2(PV/T)*delta T

Q=7/2*((1.013*10^5)*700/305)*1.7=1.38*10^6=1380 kJ

======

Part b)

Energy required, U= mgh

Mass, m= U/gh =(1.38*10^6)/(9.8*1.5)= 9.38*10^4 kg

===========